OpenStudy (anonymous):

2. Consider the numbers and . a. Compare the numbers. b. Find a rational number between the two given numbers.

OpenStudy (anonymous):

PLease help. How do I show my work with this problem on a computer?

OpenStudy (aum):

Problem is incomplete. There was probably some information provided before this problem and you need to include that.

OpenStudy (anonymous):

Oh opps yeah I thought it copied, Consider the numbers 4/13 and 3/10.

OpenStudy (anonymous):

oops*

OpenStudy (anonymous):

SO how do you solve it?

OpenStudy (aum):

4/13 and 3/10 a. Compare the numbers: To compare the numbers make the denominator the same. The least common denominator is 130. \(\Large \frac{4}{13} = \frac{4}{13} * \frac{10}{10} = \frac{40}{130} \) \(\Large \frac{3}{10} = \frac{3}{10} * \frac{13}{13} = \frac{39}{130} \) \(\Large \frac{40}{130} \gt \frac{39}{130} \) \(\Large \frac{4}{13} \gt \frac{3}{10} \)

OpenStudy (anonymous):

whoa, thanks aum :)

OpenStudy (aum):

b. Find a rational number between 4/13 and 3/10 In part a) we got 4/13 = 40/130 and 3/10 = 39/130 We need to find a rational number between 39/130 and 40/130 Multiply the top and bottom of both fractions by 2 39/130 = 78/260 40/130 = 80/260 We need a number between 78/260 and 80/260. We can choose 79/260. So 79/260 is a rational number between 4/13 and 3/10.

OpenStudy (anonymous):

Thanks a lot. Saved me :P

OpenStudy (aum):

You are welcome.

OpenStudy (anonymous):

:)